OpenStudy (anonymous):

how would verify this identity? : cos(u)cos(v)=(1/2)(cos(u-v)+cos(u+v)

OpenStudy (anonymous):

I have no idea lol.

OpenStudy (dumbcow):

using trig properties, sum and difference formulas \[\cos(u+v) = \cos u*\cos v - \sin u*\sin v\] \[\cos(u-v) = \cos u*\cos v + \sin u*\sin v\] therefore \[\cos(u-v) + \cos(u+v) = 2\cos u*\cos v\]

OpenStudy (anonymous):

oh really now

OpenStudy (dumbcow):

haha yes really